OpenStudy (anonymous):

Common brass is a copper and zinc alloy containing 37.0% zinc by mass and having a density of 8.48 g/cm3. A fitting composed of common brass has a total volume of 121.1 cm3. How many atoms (copper and zinc) does the fitting contain?

OpenStudy (aaronq):

first find the mass of the brass fitting with the volume and the density. \(\sf density=\dfrac{mass}{volume}\) then find how much mass is each, 37% is Zn and 100-37% is copper. Find the moles of each: \(\sf moles =\dfrac{mass}{Molar~mass}\) next find atoms: \(\sf moles=\dfrac{atom}{avogadro's~number}\)